OpenStudy (anonymous):

A human chain is formed by 60 people standing with their arms outstretched, each holding the hand of the person on either side. The first 30 people in the chain have arm spans of 6 feet. The next 30 people have arm spans of 5.5 feet. At the word "go!" the first person squeezes the hand of the second person. Then the second person squeezes the hand of the third, and so on. Assume that each person takes 1/3 second to pass along the signal. Which piecewise function correctly models the distance that the signal travels in t seconds?

OpenStudy (anonymous):

\[\left\{ 18t \right\} if =0 \le t \le10\] and \[\left\{ 16.5(t-10)+180 \right\} if=10<t \le 20\]

OpenStudy (anonymous):

Dang it! I just turned it in with the wrong answer. Thanks anyway. (:

OpenStudy (anonymous):

No problem, sorry I hadn't gotten here soon enough

OpenStudy (anonymous):

No, you're totally fine! I'm just happy that now I have the correct one. I can always fix it later.c: